# Armijo Line-search Can Make (Stochastic) Gradient Descent Provably Faster

### v3

@@ C.2 Helper Lemmas
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{n}{\epsilon^{2}}\right)$ implies that $f(u)\leq\frac{\epsilon^{2}}{n}$. Since $\epsilon\leq 1$ and $n\geq 1$, $f(u)\leq\epsilon$. Using this relation with the above inequality, we get,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{2n}+\frac{2\epsilon}{n}$
For $\epsilon\leq\frac{\min\left\{1,C^{\prime}\right\}}{8}\leq 1$,
$\mathbb{E}_{t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{t}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{4n}$
$\mathbb{E}[\left\|\theta_{t+1}-u\right\|_{2}^{2}]$ $\leq\mathbb{E}[\left\|\theta_{t}-u\right\|_{2}^{2}]-\frac{\min\left\{1,C^{\prime}\right\}}{4n}$
$\implies\mathbb{E}[\left\|\theta_{T}-u\right\|_{2}^{2}]$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}-\frac{\min\left\{1,C^{\prime}\right\}}{4n}\,T$ (Summing from $t=0$ to T1T-1)
$\leq\beta^{2}-\underbrace{\frac{\min\left\{1,C^{\prime}\right\}}{4n}}_{:=C}\,T$ (Since $\theta_{0}=0$ and $\left\|u\right\|=\beta$)
$T\geq\frac{\beta^{2}}{C}$ $=\frac{4\,n}{\min\{C^{\prime},1\}\,\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}$
Since the above calculations were conditioned on $G^{c}$, we can conclude that after $T=\frac{4\,n}{\min\{C^{\prime},1\}\,\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}$ iterations, $\mathbb{E}[\inf_{t\in[T]}f(\theta_{T})|G^{c}]\leq\frac{3\epsilon}{2}$. Hence, from Eq. 29
$\mathbb{E}\left[\inf_{t\in[T]}f(\theta_{t})\right]$ $\leq\frac{5\epsilon}{2}$
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,\frac{C^{\prime}}{8}\right)$ where $C^{\prime}:=\frac{1}{\lambda_{1}}=\frac{1}{648}$, SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ requires
$T=O\left(\min\left\{n,\frac{1}{\epsilon}\right\}\;\frac{1}{\gamma^{2}}\;\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$
iterations to ensure that $\mathbb{E}[\inf_{t\in[T]}f(\theta_{t})]\leq\frac{5\epsilon}{2}$.
@@
Since the algorithm guarantees that $f_{i}(\theta_{t})\geq\frac{\epsilon}{2}$ for all $t$, ${f_{t}}(\theta_{t})\geq\frac{\epsilon}{2}$. Hence,

### v4

@@ C.2 Helper Lemmas
Consider $u=\beta u^{*}$ where $\beta=\frac{1}{\gamma}\,\ln\left(\frac{n}{\epsilon^{2}}\right)$ implies that $f(u)\leq\frac{\epsilon^{2}}{n}$. Since $\epsilon\leq\frac{\min\left\{1,C^{\prime}\right\}}{8}\leq 1$ and $n\geq 1$, $f(u)\leq\epsilon$. Using this relation with the above inequality, we get,
$\mathbb{E}_{0:t}[\left\|\theta_{t+1}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t+1<\tau\}]\leq\mathbb{E}_{0:t-1}[\left\|\theta_{t}-u\right\|_{2}^{2}\,{\mathcal{I}}\{t<\tau\}]-\frac{\min\left\{1,C^{\prime}\right\}}{4n}\,\Pr[\{t<\tau\}]$
By rearranging, summing both-sides, telescoping we have,
$\frac{\min\left\{1,C^{\prime}\right\}}{4n}\,\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ $\leq\left\|\theta_{0}-u\right\|_{2}^{2}\,{\mathcal{I}}\{0<\tau\}\leq\left\|\theta_{0}-u\right\|_{2}^{2}$
$\implies\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ $\leq O\left(\frac{n}{\gamma^{2}}\,\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$ (by setting $\theta_{0}=0$ and since $\beta=\frac{1}{\gamma}\ln(\frac{n}{\epsilon^{2}})$)
Since $\tau=\sum_{t=0}^{\infty}{\mathcal{I}}\{t<\tau\}$,
$\mathbb{E}[\tau]=\mathbb{E}\left[\sum_{t=0}^{\infty}{\mathcal{I}}\{t<\tau\}\right]\leq\sum_{t=0}^{\infty}\Pr\{t<\tau\}$ (Using Fubini’s theorem)
By combining the above results we get that $\mathbb{E}[\tau]=O\left(\frac{n}{\gamma^{2}}\,\left[\ln(\frac{n}{\epsilon^{2}})\right]^{2}\right)$. ∎
@@
For logistic regression on linearly separable data with margin $\gamma$, if, for all $i$, $\left\|x_{i}\right\|\leq 1$, for a fixed $\epsilon\in\left(0,C^{\prime}\right)$ with $C^{\prime}=O(1)$, if $\tau_{\epsilon}:=\inf\{t:f(\theta_{t})\leq\epsilon\}$ is the hitting time, then SGD-SLS with $\eta_{\max}=\frac{1}{\epsilon}$ and $c=\frac{2}{3}$ ensures that,
$\mathbb{E}[\tau_{\epsilon}]=O\left(\min\left\{n,\frac{1}{\epsilon}\right\}\;\frac{1}{\gamma^{2}}\;\left[\ln\left(\frac{n}{\epsilon^{2}}\right)\right]^{2}\right)$
@@
Since the algorithm guarantees that for all $t$, ${f_{t}}(\theta_{t})\geq\frac{\epsilon}{2}$,